OpenStudy (anonymous):

Please help me out. :) Find the initial velocity needed in order to shoot a projectile vertically upward a distance of 10,000 ft.

ganeshie8 (ganeshie8):

when u shoot an object vertically up, only force acting on it is gravity

ganeshie8 (ganeshie8):

\(a = -9.8\)

ganeshie8 (ganeshie8):

u wanto use kinematics equations or calculus ?

OpenStudy (anonymous):

Calculus please. :)

OpenStudy (anonymous):

is a referring to ACCELERATION? :)

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

ahh. :)

ganeshie8 (ganeshie8):

\[a = -10 \] \[v(t)-v(0) = \int \limits_0^t a(t) dt = \int \limits_0^t-10~ dt \]

OpenStudy (anonymous):

is it integral calculus?

ganeshie8 (ganeshie8):

\[v(t)-v(0) = \int \limits_0^t a(t) dt = -10t \] \[v(t) = -10t + v(0) \]

ganeshie8 (ganeshie8):

yes, you're not familiar wid integrals yet ?

OpenStudy (anonymous):

familiar.. but only the basic.

OpenStudy (anonymous):

is it possible to use the differential calculus?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

yehey! can we use that one please? :)

ganeshie8 (ganeshie8):

seen this equation before : \(s = ut + \dfrac{1}{2}at^2\) ?

OpenStudy (anonymous):

not yet. :) just now. :)

ganeshie8 (ganeshie8):

ohk, then i would simply use kinematics equation : \(v^2 - u^2 = 2as\) at max height, velocity = 0 : \(0 - u^2 = 2as\)

ganeshie8 (ganeshie8):

plugin a = -9.8, s = 10000 and solve \(u\)

OpenStudy (loser66):

@ganeshie8 i think: \[h =h_0+V_0t +at^2\] \(h_0=0\) so that \(h= V_0t+at^2\) at the required height h = 10000, we have 10000=\(V_0t +at^2\) let it there now, \(V= V_0+1/2 at\) at the maximum height, V=0 so that \(V_0= -1/2 at^2\) replace it back to the equation above to find t, first, when we have t, replace back to find \(V_0\)

ganeshie8 (ganeshie8):

@Loser66 \(V = V_0 + at \) right ?

OpenStudy (loser66):

1/2 at, take derivative of h

ganeshie8 (ganeshie8):

what h are u talkign about ?

OpenStudy (loser66):

we know that derivative of distance is velocity, right? h = h0+ V0t +at^2 dh/dt = V = V0 +1/2at

ganeshie8 (ganeshie8):

okay, derivative of h gives u dh/dt = V0 + 2at

OpenStudy (loser66):

oh yaa. hahaha. I am sorry.

ganeshie8 (ganeshie8):

but thats incorrect still

ganeshie8 (ganeshie8):

the starting equation for h is wrong unless u define a = 2g

ganeshie8 (ganeshie8):

\[a = -g\] \[v(t) = -gt + v(0) \] \[h(t) = -g\frac{t^2}{2} + v(0)t + h(0) \]

ganeshie8 (ganeshie8):

integrating the constant acceleration backwards two times gives u above equation for height , right ?

OpenStudy (loser66):

oh, let me redo. \[h = V_0t + at^2\] \[V_0 = -2at\] \[10000= -2at^2+at^2 = -at^2 \] a =-10 so t=100 \[V_0 = -2(-10)100= 2000

ganeshie8 (ganeshie8):

doesnt look correct

OpenStudy (anonymous):

hey! im sorry. :( I've lost the connection.. :( Can you teach me again?

ganeshie8 (ganeshie8):

starting equation for h is incorrect, since you're using a = -10

ganeshie8 (ganeshie8):

and one more mistake is the given height is in feet, not meters... and the acceleration you're using is in meters/sec^2... so ..

OpenStudy (loser66):

Ok, I am sorry if it is not correct. @ganeshie8 don't bother by -10 or 10 on a . We construct the formula and work on the letter to get the last equation. After all, plug the final numbers in why do we have to plug them in too soon and then confuse anything?

ganeshie8 (ganeshie8):

Agree !

OpenStudy (loser66):

if it is just unit, so?? convert the unit, done, right?

ganeshie8 (ganeshie8):

yup !

OpenStudy (loser66):

I am sorry, I have to go. :) you help the Asker, Ok ? my friend?